# Extreme heat is associated with reductions in human activity

### v1

@@ Results
Figure 5: Implications for the future. A CMIP6 projections for days above $40^{\circ}$ in India by 2050; these indicate that Northern India could experience as many as 150 days per year at or above this temperature. B CMIP6 projects fewer days above $45^{\circ}$ but much of the North will still experience 25 days at this extreme. C and D show the consequences of this heat according to our modeling: during those days between $40-45^{\circ}$, we can expect activity to decrease by $5\%$ in rural areas and $2\%$ in urban ones; when the temperature is higher, the loss of activity could be as much as $7\%$ and $5\%$ in rural and urban areas respectively.
Figure 6: Implications for the future. A CMIP6 projections for days above $35^{\circ}$ in Indonesia by 2050; these indicate that Sumatra and Borneo could experience as many as 125 days per year at or above this temperature. B CMIP6 projects fewer days above $40^{\circ}$ but parts of Sumatra will experience 50-100 days at this level. C and D show the consequences of this heat according to our modeling: during those days between $35-40^{\circ}$, we can expect activity to decrease by $5\%$ in rural areas and $3\%$ in urban ones; when the temperature is higher, the loss of activity could be as much as $8\%$ and $6\%$ in rural and urban areas respectively.
@@
There is an important caveat when we model the entire year: warm weather during periods and in places that were once cold is good for activity, so in India our model expects aggregate mobility to fall only by 1% if we substitute the weather of 2050. Because Indonesia does not have cold weather, rising temperatures there are associated with an decline of 2% across the year. In Mexico, activity will fall as temperatures rise, by 3% on aggregate. This points to changing dynamics: we can expect heat waves to increase in both frequency and magnitude, and thus for associated disruptions to increase, but rising temperatures will have mixed effects. These effects will be spatially and temporally heterogeneous, affecting cities and seasons in different ways.
@@ Discussion
Heat waves are expected to become a regular occurrence in both rural and urban areas over the coming decades. While efforts to mitigate climate change are important, we must also develop adaptive strategies. Understanding the relationship between human behavior and extreme weather, such as heat waves, is essential for this. Our study considerably expands to a growing body of research by examining how extreme heat affects human mobility in three large developing countries. We findings show that heat waves significantly change both urban and rural mobility patterns. By examining three diverse counties across different climate zones, our research allows for comparisons and helps identify broader behavioral trends. Generally, we observed similar patterns in each context: the highest activity levels and longest trips occur at the mean temperatures for each region. However, the specifics vary between these counties. For instance, in India, activity levels were more affected by extreme high temperatures than by lower ones, whereas in Mexico and Indonesia, the reduction in activities was almost symmetric between high and low temperature extremes.
Because none of our modeling techniques are perfect, we triangulate our results with separate approaches—comparing cities before, during and after heats waves, along with additive and autoregressive models. These all point in the same direction: a nonlinear relationship between heat and activity, often with strong negative effects at the extremes. Less certain are our estimates for local differences that suggest adaptation, as average temperature correlates with optimal temperature along these nonlinear curves, but additive and autoregressive models both point to this effect. Because the relationship is noisy, further research could test the whether long-run averages are more predictive of optimal temperature than the short-run, single-year averages we compute in this study.
Our results align with existing research on the economic impacts of high temperatures [6, 25, 26] which found that productivity losses are the result of lower crop yields or impaired worker performance; we expand on this by showing that extreme heat also influences travel decisions. For example, anecdotal evidence suggests that farm workers avoid outdoor labor when temperatures exceed $40^{\circ}$ [49], and our data confirms this while pointing to broader changes in both urban and rural areas.
In India and Indonesia, respectively, we observe a noticeable decline in activity levels when average temperatures exceed 33C and 35C. We use rolling averages to show that, in India, while longer heat waves have stronger effects, activity resumes when high temperatures persist for a week or more. While extreme heat never shows a positive effect on mobility in our data, the negative effect converges to 0 when the temperature averages more than $40^{\circ}$ for 6 or 7 days. We are unable to examine the mechanisms by which this occurs but if it is the result of individuals needing to return to work, it adds to the importance of cooling efforts.
@@
A key open question remaining after our study is whether or not populations will gradually adapt with rising temperatures. This has relevance to our results because some places will have stronger heat waves, like Mexico’s arid North, and others will have higher but steadier temperatures, like its tropical South. Recent work on economic growth and public health looking across 5 decades suggests the answer is no: for most variables, responses to extreme heat have not changed [50]. An important finding in our work is that more deprived areas see stronger effects than less deprived ones, which has implications for adaptation: these areas will be amongst the last to adopt air conditioning. In the coming decades, our results may hold for vulnerable populations even if others are able to adapt.
@@ Methods
We use origin-destination matrices for three nations—India, Indonesia and Mexico—spanning 2019. These data contain flows between tessellated units, called geohashes, which are on average $\sim 22\text{km}^{2}$. Aggregated data were provided by Spectus Social Impact as part of the Netmob 2024 conference. Data are collected with the informed consent of anonymous users who have opted in to anonymized data collection for research purposes. In order to further preserve the privacy of users, all data have been aggregated by the data provider to the geohash5 level and do not include any individual-level data records.
For each geohash5 in the data, using Google Earth Engine [51] we also gather 365 days of weather data from the ERA5 daily climate aggregates [42]. These data gives us temperature, precipitation and solar radiation per day per geohash5, and it allows us to observe differences within and between regions and does not rely on weather station data. We acquire modeled estimates of these variables for each day of 2050 using spatially disaggregated data from the Coupled Model Intercomparison Project [47], an international collaboration to produce estimates of Earth’s climate through 2100. These datasets both resolve to $\sim 775\text{km}^{2}$, and thus also capture spatial heterogeneity in weather conditions. Finally, we attach data—via zonal statistics—on population [45] and deprivation [46] to each geohash5, using gridded measures estimated from remote sensing.
@@
We begin using a generalized additive model (GAM) to measure the partial effect of temperature at different levels [52]. We build a directed acyclic graph (DAG) to determine what controls are necessary in our model, shown in Fig. 2A. In addition to temperature, solar radiation and precipitation could also modulate activity, with rain dampening activity and sun heightening, so we control for these along with day-of-year. Day-of-year is modeled with a cyclic cubic spline, which captures seasonality by allowing for variation by day while forcing the value of the spline at the start to equal its value at the end. Because geography and area characteristics will also influence activity, we use geohash5 fixed effects, allowing the intercept to vary according to the unique activity profile of each geographic unit. Together, these spatial and temporal controls constitute the minimally sufficient adjustment set. While they are not confounds, adjusting for day-of-week and holiday effects improves the precision of our estimates. The resulting equation is
@@
where $T_{t}$ represents economic activity measured as Trips at time $t$, $c$ is a constant term, $\phi_{1}$ and $\phi_{7}$ are autoregressive coefficients, $\beta_{1}$ and $\beta_{2}$ are coefficients for linear and quadratic temperature effects, $\text{Temp}_{t}$ is the temperature at time $t$, $\theta_{1}$ is the moving average coefficient, and $\varepsilon_{t}$ is the error term. This model structure allows us to simultaneously account for time-dependent patterns in economic activity and the influence of temperature, providing a comprehensive framework for analyzing the complex relationships within our data.
@@ Acknowledgements
A.R. Conceptualization, methodology, investigation, writing, reviewing, editing; O.H. Conceptualization, methodology, investigation, writing, reviewing, editing.
@@
Mobility data are available as part of the NetMob 2024 Data Challenge. CMIP6 projections are accessible via Google Earth Engine. All code is available upon request.

### v2

@@ Results
Figure 5: Future implications for India. A CMIP6 projections for days above $40^{\circ}$ in India by 2050; these indicate that Northern India could experience as many as 150 days per year at or above this temperature. B CMIP6 projects fewer days above $45^{\circ}$ but much of the North will still experience 25 days at this extreme. C and D show the consequences of this heat according to our modeling: during those days between $40-45^{\circ}$, we can expect activity to decrease by $5\%$ in small cities and $2\%$ in large agglomerations; when the temperature is higher, the loss of activity could be as much as $7\%$ and $5\%$ in small and large urban areas respectively.
Figure 6: Future implications for Mexico. A CMIP6 projections for days above a given threshold in Mexico by 2050, A and B show frequency of days above $35^{\circ}$ and $40^{\circ}$, respectively; we see that the tropical South will have more frequent days above $35^{\circ}$ but the arid North will have more days above $40^{\circ}$, suggesting that heave ways will be a problem there. C and D show the consequences of these temperatures according to our model, with strong effects at higher temperatures but Mexico City, the capital and largest city, is spared of the worst effects because of its temperate climate.
@@
A key caveat in year-round modeling: warming formerly cold periods can boost activity, so in India substituting 2050 weather yields only a 1% drop overall. Because Indonesia does not have cold weather, rising temperatures there are associated with an decline of 2% across the year. In Mexico, activity will fall as temperatures rise, by 3% on aggregate. This points to changing dynamics: we can expect heat waves to increase in both frequency and magnitude, and thus for associated disruptions to increase, but rising temperatures will have mixed effects. These effects will be spatially and temporally heterogeneous, affecting cities and seasons in different ways.
@@ Discussion
Heat waves are expected to become a regular occurrence in urban areas over the coming decades. While efforts to mitigate climate change are important, we must also develop adaptive strategies. Understanding the relationship between human behavior and extreme weather, such as heat waves, is essential for this. Our study considerably expands to a growing body of research by examining how extreme heat affects human mobility in three large developing countries. We findings show that heat waves significantly change urban mobility patterns. By examining three diverse counties across different climate zones, our research allows for comparisons and helps identify broader behavioral trends. The highest activity levels and longest trips occur at the mean temperatures for each region. However, the specifics vary between these counties. For instance, in India, activity levels were more affected by extreme high temperatures than by lower ones, whereas in Mexico and Indonesia, the reduction in activities was almost symmetric between high and low temperature extremes.
Our results align with existing research on the economic impacts of high temperatures [6, 25, 26] which found that productivity losses are the result of lower crop yields or impaired worker performance; we expand on this by showing that extreme heat also influences travel decisions. For example, anecdotal evidence suggests that farm workers avoid outdoor labor when temperatures exceed $40^{\circ}$ and even service workers need to change transportation modes to reduce heat stress from walking [53], and our data confirms this while pointing to broader changes cities. Research on congestion indicates that people are willing and able to shift trips away from congested periods in the presence of taxation [54]; in showing that people substitute afternoon activity for morning activity, we find that heat functions as a tax on activity that people can avoid.
In India and Mexico, respectively, we observe a noticeable decline in activity levels when average temperatures exceed 33UTCI and 35UTCI. We use rolling averages to show that, in India, while longer heat waves have stronger effects, activity resumes when high temperatures persist for a week or more. While extreme heat never shows a positive effect on mobility in our data, the negative effect converges to 0 when the temperature averages more than $40^{\circ}$ for 6 or 7 days. We are unable to examine the mechanisms by which this occurs but if it is the result of individuals needing to return to work, it adds to the importance of cooling efforts.
@@
A limitation in our study is the coverage of the mobile phone data that we use. The data are sparser in India than they are in Mexico, with Indonesia in between, and compensate for potential biases by filtering our data to cities, where coverage is best. In Mexico, where we are most confident in the data, we see the strongest effects and the cleanest differentiation across the temperature gradient in both the GAM and TWFE models, lending credibility to our estimates. That the data in each country suggest an optimal temperature is also encouraging, but the future work will need to triangulate these findings.
@@
An remaining question is whether or not populations will gradually adapt with rising temperatures. Some places will have stronger heat waves, like Mexico’s arid North, and others will have higher but stabler temperatures, like its tropical South. Recent work on economic growth and public health looking across 5 decades suggests the answer is no: for most variables, responses to extreme heat have not changed [55]. An important finding in our work is that more deprived areas see stronger effects than less deprived ones, which has implications for adaptation: these areas will be amongst the last to adopt air conditioning. In the coming decades, our results may hold for vulnerable populations even if others are able to adapt.
@@ Methods
We use origin-destination matrices for three nations—India, Indonesia and Mexico—spanning 2019. These data contain flows between tessellated units, called geohashes. The geohashing system creates a nested hierarchy wherein longer codes indicate sharper resolution. To preserve the anonymity of users, all data have been aggregated by the data provider to the geohash5 level—with cells that are on average $\sim 22\text{km}^{2}$—and do not include any individual records. Flows are summed from trips between any geohash6 cell where a device has recurring GPS signals [56], and can thus occur in the data within the same origin and destination geohash5 as a loop. Aggregated data were provided by Spectus Social Impact as part of the Netmob 2024 conference. Data are collected with the informed consent of anonymous users who have opted in to anonymized data collection for research purposes.
For each geohash5 in the data, using Google Earth Engine [57] we also gather 365 days of weather data from the ERA5 daily climate aggregates [46]. These data gives us temperature, in the form of Universal Thermal Climate Index (UTCI), per day per geohash5, and it allows us to observe differences within and between regions without relying on weather station data. We acquire modeled estimates of these variables for each day of 2050 using spatially disaggregated data from the Coupled Model Intercomparison Project [51], an international collaboration to produce estimates of Earth’s climate through 2100. These datasets both resolve to $\sim 775\text{km}^{2}$, and thus also capture spatial heterogeneity in weather conditions. Finally, we attach data—via zonal statistics—on population [50] and deprivation [49] to each geohash5, using gridded measures estimated from remote sensing.
Using these data, we perform further checks on these data to understand is coverage and reliability in S2; while the number of devices in a given cell shows a strong correlation with the number of residents in it, we also see that more deprived areas are less represented in the data. Most of this bias comes from cells with very low population, which are also typically poor, so we clip our data to “functional urban areas” [44]. (We use multiple filters, show in LABEL:T2, to stress test our models, but in the main analysis we use the “low” constraint filter.)
Modeling the causal effect
We employ a two-way fixed effects (TWFE) approach to model the relationship between temperature anomalies and mobility in a given geohash5 cell, specified as
$\ln(y_{it})=\beta_{0}+\beta_{1}\Theta_{it}+\varphi_{i}+\nu_{t}+\varepsilon_{it}$ (1)
where $y_{it}$ represents mobility (measured as the number of trips ending in cell $i$ on day $t$). The independent variable $\Theta_{it}$ captures temperature stress in cell $i$ on day $t$, which we define as the deviation from the cell’s mean temperature over the study period. This specification allows us to examine how departures from typical local conditions affect mobility patterns, rather than absolute temperature levels which could conflate adaptation effects. The model includes two types of fixed effects: $\varphi_{i}$ represents cell-specific fixed effects that control for time-invariant characteristics of each geographic unit, such as infrastructure, elevation, or proximity to water bodies. $\nu_{t}$ captures day fixed effects that account for temporal factors affecting all cells simultaneously, including holidays, weekends, and seasonal patterns. These day fixed effects also help control for potential sampling variations in the GPS data collection, as the number of active devices may fluctuate based on the set of applications providing data on any given day.
To account for potential heterogeneity in temperature effects across different times of day, we estimate separate models for morning (6:00-12:00), afternoon (12:00-18:00), and evening (18:00-24:00) periods. This temporal disaggregation reveals how mobility responses to heat stress vary throughout the day, capturing potential behavioral adaptations such as shifting activities to cooler hours.
We cluster standard errors at the cell level to account for potential serial correlation in mobility patterns within geographic units. The coefficient of interest, $\beta_{1}$, represents the semi-elasticity of mobility with respect to temperature stress, interpretable as the percentage change in mobility associated with a one-unit increase in temperature deviation from the local mean.
@@
We begin using a generalized additive model (GAM) to measure the partial effect of temperature at different levels [58]. We build a directed acyclic graph (DAG) to determine what controls are necessary in our model, shown in Fig. S3A. In addition to temperature, solar radiation and precipitation could also modulate activity, with rain dampening activity and sun heightening, so we control for these along with day-of-year. Day-of-year is modeled with a cyclic cubic spline, which captures seasonality by allowing for variation by day while forcing the value of the spline at the start to equal its value at the end. Because geography and area characteristics will also influence activity, we use various geohash fixed effects, allowing the intercept to vary according to the unique activity profile of each geographic unit. A key assumption with this approach is that the response curve is the same across space—only the intercept will vary. Together, these spatial and temporal controls constitute the minimally sufficient adjustment set. While they are not confounds, adjusting for day-of-week and holiday effects improves the precision of our estimates. The resulting equation is
@@
where $T_{t}$ represents economic activity measured as Trips at time $t$, $UTCI$ is a constant term, $\phi_{1}$ and $\phi_{7}$ are autoregressive coefficients, $\beta_{1}$ and $\beta_{2}$ are coefficients for linear and quadratic temperature effects, $\text{Temp}_{t}$ is the temperature at time $t$, $\theta_{1}$ is the moving average coefficient, and $\varepsilon_{t}$ is the error term. This model structure allows us to simultaneously account for time-dependent patterns in economic activity and the influence of temperature, providing a comprehensive framework for analyzing the complex relationships within our data.
@@ Acknowledgements
A.R. Conceptualization, methodology, investigation, writing, reviewing, editing; O.H. Conceptualization, methodology, investigation, writing, reviewing, editing; P.B. Conceptualization, methodology, investigation, writing, reviewing, editing.
@@
Mobility data are available on request as part of the NetMob 2024 Data Challenge [56]. ERA5 climate recordings and CMIP6 climate projections are accessible via Google Earth Engine. All code is available upon request.